At what points does the graph of f(x)=x^2+3x-18 intersect the x-axis?
Answer by venugopalramana(3286)

At what points does the graph of f(x)=x^2+3x-18 intersect the x-axis?
INTERSECTION WITH X AXIS IS FOUND BY EQUATING F(X) TO ZERO AND FINDING X.
f(x)=x^2+3x-18=0
X^2+6X-3X-18=0
X(X+6)-3(X+6)=0
(X-3)(X+6)=0
X=3...OR...X=-6